OpenStudy (anonymous):

Jasmine drew a rectangle with the following properties. • The area is 8 square centimeters. • The length is twice the width. What is the perimeter of Jasmine's rectangle?

Vocaloid (vocaloid):

let L = length and W = width LW = 8 L = 2W know where to go from here?

OpenStudy (anonymous):

8X2?

Vocaloid (vocaloid):

where is X coming from?

OpenStudy (anonymous):

not sure?...

OpenStudy (anonymous):

16?

Vocaloid (vocaloid):

we have two equations LW = 8 L = 2W we can do a substitution for L in the first equation get 2W^2 = 8 solve for W

OpenStudy (anonymous):

2

OpenStudy (anonymous):

2(2)^2=8

Vocaloid (vocaloid):

good, now find L

OpenStudy (anonymous):

Let w be the width. Area is w times length or w times 2w. Solve w(2w) = 8 for w. The perimeter is 2 times the half perimeter or 2(w+2w).

OpenStudy (anonymous):

48?

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

or 128?

Vocaloid (vocaloid):

you found W, which is 2 L = 2W = ?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

What was your solution for w in the equation w(2w) = 8 or rewritten, 2 w^2 = 8 ?

Vocaloid (vocaloid):

nope

Vocaloid (vocaloid):

you width is 2 the length is twice the width, so the length is...?

OpenStudy (anonymous):

they gave no equation

OpenStudy (anonymous):

4 is length

Vocaloid (vocaloid):

good, now you have length and width now find the perimeter.

OpenStudy (anonymous):

width * length = 8 from the problem text.

OpenStudy (anonymous):

16 is perimeter

Vocaloid (vocaloid):

not quite

Vocaloid (vocaloid):

perimeter = 2(length+width) = ?

OpenStudy (anonymous):

perimeter is 8

Vocaloid (vocaloid):

again, not quite

OpenStudy (anonymous):

The perimeter is 12. Can you work backwards to find the length.

Vocaloid (vocaloid):

length + width = ?

OpenStudy (anonymous):

no this is hard for me

OpenStudy (anonymous):

12

OpenStudy (anonymous):

12 is perimeter

OpenStudy (anonymous):

length * width = area 2( length + width ) = perimeter

OpenStudy (anonymous):

@caitlynmiller 12 is correct.

OpenStudy (anonymous):

thank you